OpenStudy (anonymous):

How to find the x-intercept of f(x)=3x^2+6x+2

OpenStudy (anonymous):

0 = 3x^2 + 6x + 2 quadratic formula : (-6 +/- ((36-24)^0.5) / 6 -1 +/- (8^0.5)/6 x1,2 = -1 +/- (2^0.5)/3 so x1 = -1 + (2^0.5)/3 x2 = -1 - (2^0.5)/3